# On a class of thin obstacle-type problems for the bi-Laplacian operator

### v2

@@ 8. The free boundary: regularity and structural properties
Observe that since $\mu\in\mathbb{N}$, we must have that $u\equiv v\equiv 0$ on $S_{\mu}(u)$. We begin by proving that $d_{\mu}^{X_{0}}\leq n-1$. To this end, arguing by contradiction, assume that $d_{\mu}^{X_{0}}=n$. In particular, this assumption implies that $\nabla_{x}^{\prime}\tilde{u}_{\mu}^{X_{0}}$ and $\nabla_{x}^{\prime}\tilde{v}_{\mu}^{X_{0}}$ vanish identically on $\mathbb{R}^{n}\times\{0\}$. Consequently, the polynomials $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ must also vanish identically on $\mathbb{R}^{n}\times\{0\}$. Applying Lemma 2.1.2 in [30], we deduce that $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ are identically zero in $\mathbb{R}^{n+1}$, which contradicts ˜7.6.
@@
for all $X,X_{0}\in S_{\mu}(u)$. Assume first that $d_{\mu}^{X_{0}}=d_{u,\mu}^{X_{0}}$ and observe that, if this is the case, then $\tilde{u}_{\mu}^{X_{0}}$ is necessarily non-trivial. The desired result then readily follows by arguing as in the proof of Theorem 1.3.8 in [30]. On the other hand, if $d_{\mu}^{X_{0}}=d_{v,\mu}^{X_{0}}$, then can repeat the same argument for the non-trivial polynomial $\tilde{v}_{\mu}^{X_{0}}$. This concludes the proof. ∎
@@
    It is worth noting that the structure of the set $S_{1}(u)$ for the fourth-order problem under consideration may differ significantly from its second-order counterpart, where points with frequency $\mu=1$ are all regular. Indeed, in our case, we must still consider the scenario where $\mu=1$ and $\nabla_{x}^{\prime}u(X_{0})=0$, thereby proving that $\tilde{u}_{\mu}^{X_{0}}\equiv 0$. This, in turn, implies that $\tilde{v}_{\mu}^{X_{0}}$ cannot vanish identically. Since $\tilde{v}_{\mu}^{X_{0}}$ is independent of $y$ due to its regularity, we can find $a\in\mathbb{R}^{n}\setminus\{0\}$ such that
    $\tilde{v}_{\mu}^{X_{0}}(X)=a\cdot x.$

### v3

@@ 8. The free boundary: regularity and structural properties
Observe that since $\mu\in\mathbb{N}$, we must have that $u\equiv v\equiv 0$ on $S_{\mu}(u)$. We begin by proving that $d_{\mu}^{X_{0}}\leq n-1$. To this end, arguing by contradiction, assume that $d_{\mu}^{X_{0}}=n$. In particular, this assumption implies that $\nabla_{x}^{\prime}\tilde{u}_{\mu}^{X_{0}}$ and $\nabla_{x}^{\prime}\tilde{v}_{\mu}^{X_{0}}$ vanish identically on $\mathbb{R}^{n}\times\{0\}$. Consequently, the $\mathcal{L}_{b}$-harmonic functions $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ must also vanish identically on $\mathbb{R}^{n}\times\{0\}$. Applying Proposition 2.2 in [39], we deduce that $\tilde{u}_{\mu}^{X_{0}}$ and $\tilde{v}_{\mu}^{X_{0}}$ are identically zero in $\mathbb{R}^{n+1}$, which contradicts ˜7.6.
@@
for all $X,X_{0}\in S_{\mu}(u)$. Assume first that $d_{\mu}^{X_{0}}=d_{u,\mu}^{X_{0}}$ and observe that, if this is the case, then $\tilde{u}_{\mu}^{X_{0}}$ is necessarily non-trivial. The desired result then readily follows by arguing as in the proof of Theorem 1.3.8 in [29]. On the other hand, if $d_{\mu}^{X_{0}}=d_{v,\mu}^{X_{0}}$, then can repeat the same argument for the non-trivial function $\tilde{v}_{\mu}^{X_{0}}$. This concludes the proof. ∎
@@
    It is worth noting that the structure of the set $S_{1}(u)$ for the fourth-order problem under consideration may differ significantly from its second-order counterpart, where points with frequency $\mu=1$ are all regular. Indeed, in our case, we must still consider the scenario where $\mu=1$ and $\nabla_{x}^{\prime}u(X_{0})=0$, thereby proving that $\tilde{u}_{\mu}^{X_{0}}\equiv 0$. This, in turn, implies that $\tilde{v}_{\mu}^{X_{0}}$ cannot vanish identically. We notice also that $\tilde{v}_{\mu}^{X_{0}}$ is independent of $y$ as a consequence of the Neumann boundary condition on $B_{1}^{\prime}$ when $b\leq 0$.
@@
9. Conflict of interest and data availability statements
On behalf of all authors, the corresponding author declares that there is no conflict of interest. This manuscript has no associated data.
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewer whose remarks and suggestions substantially improved the quality of this manuscript.